# Characterizing the nuclear models informed by PREX and CREX: a view from Bayesian inference

### v1

@@ I.6 Posterior
In Fig. 8, we present the corresponding $L$ values of our sampled points of ${\Delta F^{\mathrm{Ca48}},\Delta F^{\mathrm{Pb208}}}$. As expected, $L$ is sensitive to both the neutron skin of 48Ca and 208Pb, with the points moving collectively from lower left to upper right with increasing $L$. This movement highlights the challenge of finding parameterizations that satisfy both PREX and CREX measurements by only tuning parameters controlling the $L$ values. Notably, for the same $L$, Skyrme models predict slightly smaller $\Delta F$ due to the larger slope of the constant $L$ relation for $\Delta F$ compared to RMF models, as seen in Fig. 8. This can be understood from the Pearson correlation coefficient between $\Delta F$ and $L(n_{B})$ in Fig. 9. The Pearson correlation peaks around two-thirds of the saturation density for 208Pb, where the neutron skin correlates well with the form factor difference $\Delta F$ for 208Pb. The neutron skin is known to correlate best with the symmetry energy slope at two-thirds of the saturation density for the Skyrme model Brown (2000), a correlation confirmed by RMF models as well Typel and Brown (2001). Our Pearson correlation study for $\Delta F^{\mathrm{Ca48}}$ suggests that $\Delta F^{\mathrm{Ca48}}$ measured in CREX is most sensitive to the symmetry energy slope $L$ at higher densities: $n_{B}=0.12$ fm-3 for the RMF model and $n_{B}=0.16$ fm-3 for the Skyrme model, around saturation density. The Pearson correlation is equal between the two nuclei for the RMF model, while the Pearson correlation for 208Pb is much higher than for 48Ca in Skyrme models, explaining the larger slope of the constant $L$ relation for $\Delta F$.
@@
Figure 10: Vector densities of proton (lower) and neutron (upper) in 48Ca, 90Zr and 208Pb. Different colored band represent 18%, 50% and 82% percentile under various constraints.
@@
Figure 11: Similar to Fig. 10. Different colored band represent RMF models with various $b^{\prime}_{4}$ defined in Eq.(37).
I.8 Nucleon density distribution of 48Ca and 208Pb in RMF model
@@
As highlighted in our letter, the degeneracy (correlation) between $\Delta R_{np}$ (or $\Delta F$) of 48Ca and 208Pb may be mitigated by breaking $S_{\mathbf{V}}$-$L$ relations or adjusting $b^{\prime}_{4}$. We categorized the sample into five bins based on $b^{\prime}_{4}$ and assessed the 18%, 50%, and 82% percentiles of $n(r)$ accordingly. Figure 11 elucidates how varying $b^{\prime}_{4}$ modulates the density distribution of protons and neutrons. For 48Ca, increasing $b^{\prime}_{4}$ amplifies the neutron bulge and suppresses the proton bulge around 2.5 fm. Given that the bulge around 2.5 fm exerts the greatest influence on the overall nucleon distribution due to the $r^{2}$ factor in the volume weighting, the heightened isovector density in the bulge precipitates a notable decrease in $\Delta R{np}$ and $\Delta F$ for 48Ca. The behavior of the intermediate nucleus 90Zr appears to mirror that of 48Ca, albeit with slightly smaller variations in magnitude. However, for 208Pb, increasing $b^{\prime}_{4}$ deepens the neutron deficit and enhances the proton bulge at the center, while also diminishing the isovector density at 3 fm. Although neutron excesses increase around 1.5 fm and 4 fm, the overall neutron excess remains relatively unchanged. Notably, $\Delta R_{np}$ and $\Delta F$ exhibit insensitivity or slight positive correlation with $b^{\prime}_{4}$. This can also been seen from Fig. 5.
@@
In summary, the CREX experiment necessitates a small $\Delta R_{np}$ and $\Delta F$ for 48Ca, while the PREX experiment demands a large $\Delta R_{np}$ and $\Delta F$ for 208Pb. Given the larger error bars in $\Delta F$ for the PREX experiment (three times standard deviation) compared to the CREX experiment, the joint PREX and CREX results are predominantly influenced by the CREX experiment, thereby shifting $\Delta R_{np}$ and $\Delta F$ for both 48Ca and 208Pb towards smaller values, consistent with other Bayesian analysesZhang and Chen (2023); Salinas and Piekarewicz (2023). If future experiments, such as MREX, confirm the large $\Delta F^{Pb208}$ by narrowing down the statistical error, one potential solution involves introducing a significant isospin-dependent term of the spin-orbit interaction to amplify the variation of isovector density. Such variation impacts the average interior neutron excess differently between 48Ca and 208Pb. In this study, we achieve this by adjusting $b^{\prime}_{4}$, although other interactions such as the isovector tensor interaction, may serve a similar purpose Salinas and Piekarewicz (2024). Accounting for multiple interactions that excite isovector density oscillations may offer insights into resolving the tension between PREX and CREX.

### v2

@@ Acknowledgements.
In Fig. 9, we present the corresponding $L$ values of our sampled points of $\Delta F^{\mathrm{Ca48}},\Delta F^{\mathrm{Pb208}}$}. As expected, $L$ is sensitive to both the neutron skin of 48Ca and 208Pb, with the points moving collectively from lower left to upper right with increasing $L$. This movement highlights the challenge of finding parameterizations that satisfy both PREX-2/CREX measurements by only tuning parameters controlling the $L$ values. Notably, for the same $L$, Skyrme models predict slightly smaller $\Delta F$ due to the larger slope of the constant $L$ relation for $\Delta F$ compared to RMF models, as seen in Fig. 9. This can be understood from the Pearson correlation coefficient between $\Delta F$ and $L(n_{B})$ in Fig. 11.
In Fig. 10, we present a more comprehensive demonstration of the $b^{\prime}_{4}$ correlation with $\Delta F^{\mathrm{Ca48}},\Delta F^{\mathrm{Pb208}}$} comparing with existing Skyrme and RMF models. The widely used SLy4 model has $W_{0}=2b_{4}=2b_{4}^{\prime}=123$ MeV fm5Chabanat et al. (1998a), equivalent to $b_{4}=b^{\prime}_{4}=0.312$ fm4, which lies on the 2-$\sigma$ contour of the PREX-2/CREX measurement. The first Skyrme models with $b_{4}\neq b^{\prime}_{4}$, SkI3 and SkI4 are introduced with $b^{\prime}_{4}\leq 0<b_{4}$, placing them further away from the PREX-2/CREX measurement Reinhard and Flocard (1995b). SV classes of Skyrme model prefer $0<b^{\prime}_{4}/b_{4}<0.5$, see Fig. 9 in Klüpfel et al. (2009). TOV-min takes a relatively large $b^{\prime}_{4}=0.426$Erler et al. (2013), placing it closer towards the right bottom corner. More recent calibrated models UNEDF0 with $b_{4}=0.634$, $b^{\prime}_{4}=-0.463$ and UNEDF1 with $b_{4}=0.194$, $b^{\prime}_{4}=0.362$ have $b^{\prime}_{4}$ of opposite signKortelainen et al. (2012). Therefore, UNEDF1 is much closer to the PREX-2/CREX measurements compared to UNEDF0. Most recently, the extended Skyrme model with $b^{\prime}_{4}=1.27$ is proposed to reach the center of PREX-2/CREX measurementYue et al. (2024), which is consistent with our correlation which gives $b^{\prime}_{4}=1.37\pm 0.49$ fm4 for Skyrme models from PREX-2/CREX measurements of $\Delta F$. In the case of RMF models, the effective parameter is $b^{\prime}_{4}=1.02\pm 0.37$, corresponding to a 90% lower bound of $b^{\prime}_{4}>0.544$ fm4 or $g_{\delta}^{2}+1.65g_{\rho}^{2}>2433$ after incorporating the masses of nucleons and mesons. Widely used RMF models without delta mesons, such as FSUGold2 Chen and Piekarewicz (2014) with $g_{\rho}^{2}=80.4656$ and FSUGarnet Chen and Piekarewicz (2015) with $g_{\rho}^{2}=192.9274$, typically exhibit a small effective $b^{\prime}_{4}<0.1$. Other RMF models with $b^{\prime}_{4}\lesssim 0.1$, such as NL3Lalazissis et al. (1997b), IOPB-IKumar et al. (2018), IUFSUFattoyev et al. (2010), BigAppleFattoyev et al. (2020), HPUA-CSharma et al. (2023), all lie beyond the 2-$\sigma$ region, although these models span a wide range of symmetry energy slope L. New RMF models featuring $\delta$ mesons with small isovector Yukawa coupling, such as BSRV with $g_{\delta}^{2}<20$ and $g_{\rho}^{2}<265$, corresponding to $b^{\prime}_{4}\lesssim 0.1$, also do not show significant improvement Kumar et al. (2023). However, models like DINOa-c, which display large isovector Yukawa couplings ($g_{\delta}^{2}>1100$ and $g_{\rho}^{2}>800$), all surpass the 90% lower bound of $b^{\prime}_{4}>0.544$ Reed et al. (2024). Consequently, DINOa-c lies close to the 1-$\sigma$ confidence region of PREX-2/CREX.
In Fig.11, the Pearson correlation peaks around two-thirds of the saturation density for 208Pb, where the neutron skin correlates well with the form factor difference $\Delta F$ for 208Pb. The neutron skin is known to correlate best with the symmetry energy slope at two-thirds of the saturation density for the Skyrme model Brown (2000), a correlation confirmed by RMF models as well Typel and Brown (2001). Note that for both Skyrme and RMF models, the Pearson correlation between P and $\Delta F^{Pb208}$ is higher than that between P and $\Delta F^{Ca208}$ below and around the saturation density. It clearly shows the neutron skin of $^{208}\mathrm{Pb}$ is more sensitive to the pressures in neutron stars than $^{208}\mathrm{Ca}$ does. Additionally, our Pearson correlation study for $\Delta F^{\mathrm{Ca48}}$ suggests that $\Delta F^{\mathrm{Ca48}}$ measured in CREX is most sensitive to the symmetry energy slope $L$ at $n_{B}=0.12$ fm-3 for the RMF model and $n_{B}=0.16$ fm-3 for the Skyrme model, around saturation density. The Pearson correlation with L are about the same between the two nuclei in the RMF models, while the Pearson correlation with L and for 208Pb is much higher than that for 48Ca in Skyrme models, explaining the larger slope of the constant $L$ relation for $\Delta F$.
@@
Figure 12: Vector densities of protons (lower) and neutrons (upper) in 48Ca, 90Zr, and 208Pb. The different colored bands and central lines represent the 18%, 50%, and 82% percentiles under various constraints specified in the legend.
@@
Figure 13: Similar to Fig. 12. The different colored bands and central lines represent RMF models with various $b^{\prime}_{4}$ defined in Eq.(35).
.9 Nucleon density distribution of 48Ca and 208Pb in RMF model
@@
As highlighted in our letter, the degeneracy (correlation) between $\Delta R_{np}$ (or $\Delta F$) of 48Ca and 208Pb may be mitigated by breaking $S_{\mathbf{V}}$-$L$ relations or adjusting $b^{\prime}_{4}$. We categorized the sample into five bins based on $b^{\prime}_{4}$ and assessed the 18%, 50%, and 82% percentiles of $n(r)$ accordingly. Figure 13 elucidates how varying $b^{\prime}_{4}$ modulates the density distribution of protons and neutrons. For 48Ca, increasing $b^{\prime}_{4}$ amplifies the neutron bulge and suppresses the proton bulge around 2.5 fm. Given that the bulge around 2.5 fm exerts the greatest influence on the overall nucleon distribution due to the $r^{2}$ factor in the volume weighting, the heightened isovector density in the bulge precipitates a notable decrease in $\Delta R_{np}$ and $\Delta F$ for 48Ca. The behavior of the intermediate nucleus 90Zr appears to mirror that of 48Ca, albeit with slightly smaller variations in magnitude. However, for 208Pb, increasing $b^{\prime}_{4}$ deepens the neutron deficit and enhances the proton bulge at the center, while also diminishing the isovector density at 3 fm. Although neutron excesses increase around 1.5 fm and 4 fm, the overall neutron excess remains relatively unchanged. Notably, $\Delta R_{np}$ and $\Delta F$ exhibit insensitivity or slight positive correlation with $b^{\prime}_{4}$. This can also be seen from Pearson correlation of $\Delta F$ in Fig. 5 and Fig. 6.
@@
In summary, the CREX experiment necessitates a small $\Delta R_{np}$ and $\Delta F$ for 48Ca, while the PREX experiment demands a large $\Delta R_{np}$ and $\Delta F$ for 208Pb. Given the larger error bars in $\Delta F$ for the PREX experiment (three times standard deviation) compared to the CREX experiment, the joint PREX-2/CREX results are predominantly influenced by the CREX experiment, thereby shifting $\Delta R_{np}$ and $\Delta F$ for both 48Ca and 208Pb towards smaller values, consistent with other Bayesian analysesZhang and Chen (2023); Salinas and Piekarewicz (2023). If future experiments, such as MREX, confirm the large $\Delta F^{Pb208}$ by narrowing down the statistical error, one potential solution involves introducing a significant isospin-dependent term of the spin-orbit interaction to amplify the variation of isovector density. Such variation impacts the average interior neutron excess differently between 48Ca and 208Pb. In this study, we achieve this by adjusting $b^{\prime}_{4}$, although other interactions such as the isovector tensor interaction, may serve a similar purpose Salinas and Piekarewicz (2024); Yue et al. (2024). Accounting for multiple interactions that excite isovector density oscillations may offer insights into resolving the tension between PREX-2/CREX.